# LLM4Cell: A Survey of Large Language and Agentic Models for Single-Cell Biology

### v2

@@ 6.2 Batch Effects and Heterogeneity
Robustness to batch effects varies widely. scFoundation  (Hao et al., 2024), CellFM  (Zeng et al., 2025), and iSEEEK  (Shen et al., 2022) mitigate donor and protocol variation through large-scale multi-domain pretraining, while others rely on explicit batch tokens or adversarial alignment. True cross-platform generalization remains limited; even large models show degradation when tested on unseen sequencing chemistries or tissue types.
@@ 6.3 Multi-omics Alignment
Models such as scMGPT  (Palayew et al., 2025), GET  (Fu et al., 2025), and ChromFound  (Jiao et al., 2025) demonstrate consistent alignment across RNA, ATAC, and protein features using cross-attention or contrastive fusion. Spatial hybrids (HEIST, FmH2ST) further connect molecular and visual domains. While alignment quality improves, limited paired datasets and uneven modality coverage constrain reproducibility and benchmarking.
@@ 6.5 Cross-Species and Cross-Tissue Generalization
Cross-organism transfer remains challenging. iSEEEK  (Shen et al., 2022), UCE  (Rosen et al., 2023), and GeneCompass  (Yang et al., 2024b) align human and mouse cell embeddings via ortholog mapping, while scPlantLLM  (Cao et al., 2025) extends this to plants. Performance drops sharply for divergent taxa, reflecting incomplete ortholog tables and bias toward human datasets.
@@ 6.6 Atlas Fairness and Representation Balance
Dataset imbalance affects nearly all models. Human and immune-cell–dominant atlases overrepresent certain tissues and demographics. Teddy (Chevalier et al., 2025) introduces benchmarking for fairness and representation balance, but fairness metrics are rarely adopted. No model yet enforces demographic or tissue-specific parity during training.
@@ 6.7 Explainability and Interpretability
Text-bridge systems (GenePT, CellLM, Cell2Text  (Kharouiche et al., 2025)) improve interpretability by linking embeddings to ontology terms. Agentic systems (scAgent) extend this via natural-language reasoning and tool explanations. Most foundation models still rely on attention visualization, providing limited causal transparency.
@@ 6.8 Privacy and Ethics
Data privacy remains underexplored. Synthetic generation (scGFT  (Nouri, 2025)) and open-access chat systems (ChatCell  (Fang et al., 2024)) raise concerns over re-identification and data leakage. No single-cell LLM currently implements federated or privacy-preserving learning, and ethical guidance for multi-omic sharing remains informal.
@@ 6.9 Scalability and Efficiency
Scaling trends mirror NLP: efficient transformers (xTrimoGene (Gong et al., 2023)) and state-space models (GeneMamba (Qi et al., 2025a), scMamba (Yuan et al., 2025)) reduce memory cost while retaining accuracy. CellFM  (Zeng et al., 2025) demonstrates training across 100 M cells, showing feasibility of atlas-scale learning. Still, high compute requirements limit accessibility for most research groups.
@@ 6.10 Emerging Paradigms and Agentic Behavior
Recent frameworks (scAgent  (Mao et al., 2025), CellVerse  (Zhang et al., 2025b), EpiAgent (Chen et al., 2025b)) introduce reasoning and autonomous decision pipelines, integrating LLM controllers with specialized encoders. These systems achieve the highest scores in explainability and cross-modal planning but lack standardized evaluation of reasoning fidelity or reproducibility.
@@ C.1 Foundation Models (Extended Details)
Most adopt Transformer backbones with gene tokens as contextual units. tGPT and scBERT encode ranked or discrete gene sequences, while Geneformer and scFoundation use masked-token modeling. Hybrid designs such as scGraphformer  (Fan et al., 2024) and scRDiT  (Dong et al., 2025) incorporate graph or diffusion-based attention, improving spatial and regulatory context.
@@ C.5 Agentic Frameworks (Extended Details)
Agentic models couple (i) a pretrained molecular encoder, (ii) an LLM-based controller, and (iii) external tool interfaces. scAgent executes multi-step annotation workflows via ontology queries; CellVerse coordinates multiple domain agents for transcriptomic, spatial, and literature reasoning; EpiAgent extends agentic control to enhancer–gene analysis; and Teddy/scPilot (Joodaki et al., 2024) provide lightweight orchestrators for benchmarking pipelines. Grounding sources include the Human Cell Atlas ontology, UBERON, and PubMed. These systems score highest in Explainability and Emerging paradigms, marking a transition from static embeddings to interactive, goal-directed modeling. Open challenges include evaluation of reasoning accuracy, privacy, and reproducibility.
@@ Appendix E Methods Comparison
scGPT Cui et al. (2024) Nature Methods Foundation Multiomics (scRNA + optionally multi-omic modes) Atlas (trained on large atlas of single-cell datasets) No Annotation, Integration, Perturbation (annotation as primary) 6
@@
Geneformer Theodoris et al. (2023) Nature Foundation scRNA None / Rank-based (implicit) No Cell classification, in-silico perturbation, network prediction 6
@@
scFoundation Hao et al. (2024) Nature Methods Foundation scRNA / multi-omics (transcriptomics focus) Atlas (large pretrained corpus) No Cell annotation, perturbation prediction, drug response, gene module inference 7
@@
CellFM Zeng et al. (2025) Nature Communications Foundation scRNA Atlas / value-projection No Learns embeddings from 100M cells and supports annotation, perturbation, gene function 7
@@
iSEEEK Shen et al. (2022) Briefings in Bioinformatics Foundation scRNA Atlas (11.9M cells, human + mouse) No Integrates massive single-cell datasets via gene-ranking similarity to enable scalable cross-dataset embedding and clustering 5
@@
tGPT Shen et al. (2023) iScience Foundation Model (Autoregressive) scRNA Atlas (22.3 M cells) No Treats gene-expression ranks as token sequences and learns generative embeddings for cell clustering, trajectory inference, and bulk-tissue analysis 5
@@
scBERT Yang et al. (2022) Nature Machine Intelligence Foundation Model (Encoder-only) scRNA Atlas (public scRNA-seq corpora) No Learns bidirectional gene–cell embeddings using BERT-style masked modeling for robust cell-type annotation and novel cell discovery 4
@@
UCE (Universal Cell Embeddings) Rosen et al. (2023) BioRxiv Foundation scRNA (cross-species) Atlas / binary masked self-supervision No Embeds any cell zero-shot across species into a shared latent space for clustering, lineage inference, annotation 6
@@
GeneCompass Yang et al. (2024b) Nature-/Cell Research Foundation Model scRNA / cross-species Ontology (GRN + co-expression + gene family knowledge) No Learns cross-species gene–cell embeddings and supports annotation, perturbation, dose-response, and GRN inference 6
@@
scELMo Liu et al. (2023) BioRxiv Text-Bridge LLM scRNA + Metadata Text LLM-generated embeddings from gene/metadata descriptions + raw data No Embeds cells via text-derived gene metadata embeddings combined with expression, then supports clustering, batch correction, annotation, perturbation 5
@@
CellPLM Wen et al. (2023) ICLR Spatial (Multimodal Foundation Model) scRNA + Spatial (SRT) Atlas + Spatial relations (uses SRT in pretraining) No Treats cells as tokens and tissues as sentences, leveraging spatially-resolved transcriptomics and a Gaussian-mixture prior to encode inter-cell relations for denoising, spatial imputation, and perturbation prediction. 4
@@
scMoFormer Tang et al. (2023) ArXiv Spatial (Multimodal Foundation Model) scRNA + Protein / Multi-omics Atlas + domain knowledge in cross-modality aggregation No Uses modality-specific transformers and cross-attention to impute missing modalities, classify cells, and fuse multimodal representations 4
@@
scFormer Xu et al. (2024) ArXiv Spatial (Multimodal Foundation Model) (or Foundation + multi-omics) Transcript- omics + Proteomics / multimodal Atlas + transformer-based fusion No Aligns and integrates multi-omics single-cell data, recovers missing modalities, and transfers labels across modalities 4
@@
scMulan Bian et al. (2024a) BioRxiv Foundation Model scRNA + Metadata Atlas + prompt-conditioned generative modeling No Encodes each cell as a “c-sentence” integrating expression + metadata; supports zero-shot annotation, batch integration, and conditional generation 3
@@
scPRINT Kalfon et al. (2025) Nature Communications Foundation Model scRNA Atlas (50M+ cell pretraining) No Learns cell embeddings and infers cell-specific gene regulatory networks; supports zero-shot denoising, batch correction, label prediction, expression reconstruction 5
@@
scGraphformer Fan et al. (2024) Nature Communications Biology Foundation Model scRNA Atlas + relational prior (kNN bias, refined) No Learns a cell–cell graph via a transformer-GNN hybrid for better classification and interaction inference 4
@@
scRDiT Dong et al. (2025) ArXiv Foundation Model scRNA (transcriptome) Atlas-like generative prior / diffusion modeling No Generates synthetic scRNA-seq samples via diffusion transformer + DDIM for accelerated sampling 2
@@
scGFT Nouri (2025) Nature Communications Biology Foundation Model scRNA Fourier-based perturbation / reconstruction (train-free) No Synthesizes new single-cell expression profiles by perturbing Fourier components in frequency space. 4
@@
scTrans Lu et al. (2024) IJCAI Foundation Model scRNA Sub-vector masked completion over gene modules No Learns multi-scale sub-vector tokens to perform gene-selective cell-type annotation via masked completion and contrastive regularization 3
@@
scGT (Graph Transformer) Qi et al. (2025b) Bioinfo-rmatics Advance Graph / Multi-omics Integration scRNA + scATAC Observed / Hybrid Graph No Multi-omics integration + label transfer 4
@@
TransformerST Lu et al. (2024) Briefings in Bioinformatics Spatial / Multimodal FM Spatial (histology + gene expression) Spatial (histology + gene expression) No Super-resolution gene expression & tissue clustering 3
@@
scGPT-spatial Cui et al. (2024) BioRxiv Spatial (Multimodal Foundation Model) Spatial Transcriptomics + scRNA prior Atlas + spatial-aware decoding No Extends scGPT via continual pretraining to spatial data, supports multi-slide integration, cell-type deconvolution, and spatial gene imputation 3
@@
HEIST Madhu et al. (2025) BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + proteomics Hierarchical graph modeling (spatial + GRNs) No Learns joint embeddings of cells and genes in spatial context to perform cell annotation, gene imputation, spatial clustering, clinical outcome prediction 4
@@
stFormer Cao et al. (2024) BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + ligand context Atlas + biased cross-attention to ligand niche genes No Learns gene embeddings contextualized by ligand signals in spatial microenvironments; aids clustering, ligand–receptor inference, and perturbation simulation 4
@@
FmH2ST Wang et al. (2025) Nucleic Acids Research Spatial (Multimodal Foundation Model) Histology image + spatial transcriptomics Image foundation + dual graphs + spot branch No Predict spatial gene expression from histology using fused image and spot features, supporting denoising, heterogeneity detection, and regulatory inference 3
@@
OmiCLIP Cui et al. (2025) Nature Methods Spatial (Multimodal Foundation Model) Histology + Spatial Transcriptomics Image–gene contrastive alignment (rank-based) No Learns unified visual-omics embeddings linking histopathology and spatial gene expression; enables cross-modal prediction, annotation, and tissue retrieval 7
@@
QuST-LLM Huang (2024) ArXiv Text-Bridge LLM Spatial transcriptomics + histology metadata GO-term + gene enrichment + LLM narrative overlay Yes Converts ST data and ROIs into human-readable narratives and matches natural language queries to spatial regions via GO/LLM interpretation 4
@@
SpaCCC Yang et al. (2024a) IEEE Xplore Text-Bridge LLM Spatial transcriptomics / transcriptomic genes (LRs) LLM embeddings of ligand + receptor genes No Infers spatially resolved cell–cell communication by embedding LR pairs in LLM latent space + diffusion / permutation test filtering 3
@@
spaLLM Ji et al. (2024) Briefings in Bioinformatics Spatial (Multimodal Foundation Model) Spatial multi-omics (RNA, ATAC, protein) scGPT embeddings + GNN + attention fusion No Enhances spatial domain identification by fusing LLM-derived embeddings and spatial-omics signals via multi-view attention 3
@@
scMMGPT (2025) Shi et al. (2025) ArXiv Text-Bridge LLM / Multimodal_FM RNA + Text (metadata) Annotation, description No Generation - links single-cell and text PLMs to describe cells, generate pseudo-cells from text, and enhance annotation through text-conditioned reasoning 6
@@
Cell2Sentence (C2S) Levine et al. (2024) PMLR Text-Bridge LLM RNA (scRNA-seq converted to “cell sentences”) Marker / Implicit (gene rank-order) No Generation, annotation & reconstruction - encodes cells as gene-ranked sentences and fine-tunes LLMs to classify or generate biologically meaningful cell text 6
@@
C2S-Scale Rizvi et al. (2025) BioRxiv Text-Bridge LLM+ Multimodal Foundation Model RNA + Text / Metadata Atlas + Text / Implicit rank grounding No Chat, generation & annotation - trains large LLMs on cell sentences and biological text to enable perturbation prediction and multicellular summarization 7
@@
Cell2Text Kharouiche et al. (2025) ArXiv Text-Bridge LLM + Multimodal Foundation Model RNA → Natural Language Gene-level embeddings + ontology / metadata grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict expression and interpret TF–cis interactions 7
@@
GenePT Chen and Zou (2024) BioRxiv Foundation Model / Text-Augmented scRNA Literature / Text embedding grounding No Embedding & downstream prediction - uses GPT-3.5 gene text embeddings to derive cell embeddings via weighted or ranked gene aggregation. 5-7
@@
CellLM Zhao et al. (2023) ArXiv Foundation Model / Representation model RNA (single cell expression) Implicit - contrastive embedding from expression data No Represent cells via a contrastive-learning transformer, optimize embedding space for tasks like cell-type annotation, drug sensitivity prediction 6
@@
scExtract Wu and Tang (2025) Genome Biology Agentic Framework (Text-Bridge LLM) RNA (scRNA-seq) + Text Article-based parameter extraction Yes Annotation & Integration - uses LLMs to extract pipeline parameters from publications and apply them for dataset harmonization. 7
@@
scAgent Mao et al. (2025) ArXiv Agentic Framework (Text-Bridge hybrid) RNA (scRNA-seq) Reference atlas + marker gene reasoning + memory grounding Yes Universal cell annotation & novel cell discovery - scAgent uses an LLM planning module, memory, and tool modules to annotate cells across tissues, detect unknown types, and incrementally learn new annotations 7
@@
EpiFoundation Wu et al. (2025) BioRxiv Epigenomic Foundation Model / Multimodal alignment scATAC (chromatin accessibility) Peak-to-gene supervision (alignment to expression) No Cell embedding, annotation, batch correction, gene expression prediction - trains on sparse peak sets, aligns to gene expression supervision, and transfers learned embeddings for downstream ATAC tasks 7
@@
EpiAgent Chen et al. (2025b) BioRxiv Epigenomic Foundation Model / Multimodal text-bridge hybrid scATAC / chromatin accessibility Peak tokenization + external embeddings + regulatory supervision grounding No Embedding, annotation, imputation, and perturbation prediction - encodes chromatin accessibility as ranked cCRE tokens, enabling zero-shot cell annotation, peak imputation, and response prediction. 8
@@
ChromFound Jiao et al. (2025) Nature Epigenomic → Expression Transformer / Regulatory FM scATAC (chromatin accessibility) + DNA sequence Motif × peak matrix / masked regulatory grammar grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict gene expression in seen and unseen cell types; also interprets transcription factor interactions and cis-regulatory elements 8
@@
GET (General Expression Transformer) Fu et al. (2025) Nature Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6
@@
scMPT Palayew et al. (2025) ArXiv Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6

### v3

@@ 6.2 Batch Effects and Heterogeneity
Robustness to batch effects varies widely. scFoundation  (hao2024large), CellFM  (zeng2025cellfm), and iSEEEK  (shen2022universal) mitigate donor and protocol variation through large-scale multi-domain pretraining, while others rely on explicit batch tokens or adversarial alignment. True cross-platform generalization remains limited; even large models show degradation when tested on unseen sequencing chemistries or tissue types.
@@ 6.3 Multi-omics Alignment
Models such as scMGPT  (palayew2025towards), GET  (fu2025foundation), and ChromFound  (jiao2025chromfound) demonstrate consistent alignment across RNA, ATAC, and protein features using cross-attention or contrastive fusion. Spatial hybrids (HEIST, FmH2ST) further connect molecular and visual domains. While alignment quality improves, limited paired datasets and uneven modality coverage constrain reproducibility and benchmarking.
@@ 6.5 Cross-Species and Cross-Tissue Generalization
Cross-organism transfer remains challenging. iSEEEK  (shen2022universal), UCE  (rosen2023universal), and GeneCompass  (yang2024genecompass) align human and mouse cell embeddings via ortholog mapping, while scPlantLLM  (scplantlmm) extends this to plants. Performance drops sharply for divergent taxa, reflecting incomplete ortholog tables and bias toward human datasets.
@@ 6.6 Atlas Fairness and Representation Balance
Dataset imbalance affects nearly all models. Human and immune-cell–dominant atlases overrepresent certain tissues and demographics. Teddy (chevalier2025teddy) introduces benchmarking for fairness and representation balance, but fairness metrics are rarely adopted. No model yet enforces demographic or tissue-specific parity during training.
@@ 6.7 Explainability and Interpretability
Text-bridge systems (GenePT, CellLM, Cell2Text  (kharouiche2025cell2text)) improve interpretability by linking embeddings to ontology terms. Agentic systems (scAgent) extend this via natural-language reasoning and tool explanations. Most foundation models still rely on attention visualization, providing limited causal transparency.
@@ 6.8 Privacy and Ethics
Data privacy remains underexplored. Synthetic generation (scGFT  (nouri2025single)) and open-access chat systems (ChatCell  (fang2024chatcell)) raise concerns over re-identification and data leakage. No single-cell LLM currently implements federated or privacy-preserving learning, and ethical guidance for multi-omic sharing remains informal.
@@ 6.9 Scalability and Efficiency
Scaling trends mirror NLP: efficient transformers (xTrimoGene (gong2023xtrimogene)) and state-space models (GeneMamba (qi2025bidirectional), scMamba (yuan2025scmamba)) reduce memory cost while retaining accuracy. CellFM  (zeng2025cellfm) demonstrates training across 100 M cells, showing feasibility of atlas-scale learning. Still, high compute requirements limit accessibility for most research groups.
@@ 6.10 Emerging Paradigms and Agentic Behavior
Recent frameworks (scAgent  (mao2025scagent), CellVerse  (zhang2025cellverse), EpiAgent (chen2025epiagent)) introduce reasoning and autonomous decision pipelines, integrating LLM controllers with specialized encoders. These systems achieve the highest scores in explainability and cross-modal planning but lack standardized evaluation of reasoning fidelity or reproducibility.
@@ 8 Conclusions
Acknowledgments
This work was supported in part by Virginia Tech, the Department of Computer Science, and the U.S. National Science Foundation (NSF) under Awards #2125798, #2344169, and #2319522. We gratefully acknowledge their support and resources that made this research possible.
@@ C.1 Foundation Models (Extended Details)
Most adopt Transformer backbones with gene tokens as contextual units. tGPT and scBERT encode ranked or discrete gene sequences, while Geneformer and scFoundation use masked-token modeling. Hybrid designs such as scGraphformer  (fan2024scgraphformer) and scRDiT  (dong2025scrdit) incorporate graph or diffusion-based attention, improving spatial and regulatory context.
@@ C.5 Agentic Frameworks (Extended Details)
Agentic models couple (i) a pretrained molecular encoder, (ii) an LLM-based controller, and (iii) external tool interfaces. scAgent executes multi-step annotation workflows via ontology queries; CellVerse coordinates multiple domain agents for transcriptomic, spatial, and literature reasoning; EpiAgent extends agentic control to enhancer–gene analysis; and Teddy/scPilot (joodaki2024detection) provide lightweight orchestrators for benchmarking pipelines. Grounding sources include the Human Cell Atlas ontology, UBERON, and PubMed. These systems score highest in Explainability and Emerging paradigms, marking a transition from static embeddings to interactive, goal-directed modeling. Open challenges include evaluation of reasoning accuracy, privacy, and reproducibility.
@@ Appendix E Methods Comparison
scGPT cui2024scgpt Nature Methods Foundation Multiomics (scRNA + optionally multi-omic modes) Atlas (trained on large atlas of single-cell datasets) No Annotation, Integration, Perturbation (annotation as primary) 6
@@
Geneformer theodoris2023transfer Nature Foundation scRNA None / Rank-based (implicit) No Cell classification, in-silico perturbation, network prediction 6
@@
scFoundation hao2024large Nature Methods Foundation scRNA / multi-omics (transcriptomics focus) Atlas (large pretrained corpus) No Cell annotation, perturbation prediction, drug response, gene module inference 7
@@
CellFM zeng2025cellfm Nature Communications Foundation scRNA Atlas / value-projection No Learns embeddings from 100M cells and supports annotation, perturbation, gene function 7
@@
iSEEEK shen2022universal Briefings in Bioinformatics Foundation scRNA Atlas (11.9M cells, human + mouse) No Integrates massive single-cell datasets via gene-ranking similarity to enable scalable cross-dataset embedding and clustering 5
@@
tGPT shen2023generative iScience Foundation Model (Autoregressive) scRNA Atlas (22.3 M cells) No Treats gene-expression ranks as token sequences and learns generative embeddings for cell clustering, trajectory inference, and bulk-tissue analysis 5
@@
scBERT yang2022scbert Nature Machine Intelligence Foundation Model (Encoder-only) scRNA Atlas (public scRNA-seq corpora) No Learns bidirectional gene–cell embeddings using BERT-style masked modeling for robust cell-type annotation and novel cell discovery 4
@@
UCE (Universal Cell Embeddings) rosen2023universal BioRxiv Foundation scRNA (cross-species) Atlas / binary masked self-supervision No Embeds any cell zero-shot across species into a shared latent space for clustering, lineage inference, annotation 6
@@
GeneCompass yang2024genecompass Nature-/Cell Research Foundation Model scRNA / cross-species Ontology (GRN + co-expression + gene family knowledge) No Learns cross-species gene–cell embeddings and supports annotation, perturbation, dose-response, and GRN inference 6
@@
scELMo liu2023scelmo BioRxiv Text-Bridge LLM scRNA + Metadata Text LLM-generated embeddings from gene/metadata descriptions + raw data No Embeds cells via text-derived gene metadata embeddings combined with expression, then supports clustering, batch correction, annotation, perturbation 5
@@
CellPLM wen2023cellplm ICLR Spatial (Multimodal Foundation Model) scRNA + Spatial (SRT) Atlas + Spatial relations (uses SRT in pretraining) No Treats cells as tokens and tissues as sentences, leveraging spatially-resolved transcriptomics and a Gaussian-mixture prior to encode inter-cell relations for denoising, spatial imputation, and perturbation prediction. 4
@@
scMoFormer tang2023single ArXiv Spatial (Multimodal Foundation Model) scRNA + Protein / Multi-omics Atlas + domain knowledge in cross-modality aggregation No Uses modality-specific transformers and cross-attention to impute missing modalities, classify cells, and fuse multimodal representations 4
@@
scFormer xu2024scmformer ArXiv Spatial (Multimodal Foundation Model) (or Foundation + multi-omics) Transcript- omics + Proteomics / multimodal Atlas + transformer-based fusion No Aligns and integrates multi-omics single-cell data, recovers missing modalities, and transfers labels across modalities 4
@@
scMulan bian2024scmulan BioRxiv Foundation Model scRNA + Metadata Atlas + prompt-conditioned generative modeling No Encodes each cell as a “c-sentence” integrating expression + metadata; supports zero-shot annotation, batch integration, and conditional generation 3
@@
scPRINT kalfon2025scprint Nature Communications Foundation Model scRNA Atlas (50M+ cell pretraining) No Learns cell embeddings and infers cell-specific gene regulatory networks; supports zero-shot denoising, batch correction, label prediction, expression reconstruction 5
@@
scGraphformer fan2024scgraphformer Nature Communications Biology Foundation Model scRNA Atlas + relational prior (kNN bias, refined) No Learns a cell–cell graph via a transformer-GNN hybrid for better classification and interaction inference 4
@@
scRDiT dong2025scrdit ArXiv Foundation Model scRNA (transcriptome) Atlas-like generative prior / diffusion modeling No Generates synthetic scRNA-seq samples via diffusion transformer + DDIM for accelerated sampling 2
@@
scGFT nouri2025single Nature Communications Biology Foundation Model scRNA Fourier-based perturbation / reconstruction (train-free) No Synthesizes new single-cell expression profiles by perturbing Fourier components in frequency space. 4
@@
scTrans lu2024sctrans IJCAI Foundation Model scRNA Sub-vector masked completion over gene modules No Learns multi-scale sub-vector tokens to perform gene-selective cell-type annotation via masked completion and contrastive regularization 3
@@
scGT (Graph Transformer) qi2025scgt Bioinfo-rmatics Advance Graph / Multi-omics Integration scRNA + scATAC Observed / Hybrid Graph No Multi-omics integration + label transfer 4
@@
TransformerST lu2024sctrans Briefings in Bioinformatics Spatial / Multimodal FM Spatial (histology + gene expression) Spatial (histology + gene expression) No Super-resolution gene expression & tissue clustering 3
@@
scGPT-spatial cui2024scgpt BioRxiv Spatial (Multimodal Foundation Model) Spatial Transcriptomics + scRNA prior Atlas + spatial-aware decoding No Extends scGPT via continual pretraining to spatial data, supports multi-slide integration, cell-type deconvolution, and spatial gene imputation 3
@@
HEIST madhu2025heist BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + proteomics Hierarchical graph modeling (spatial + GRNs) No Learns joint embeddings of cells and genes in spatial context to perform cell annotation, gene imputation, spatial clustering, clinical outcome prediction 4
@@
stFormer cao2024stformer BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + ligand context Atlas + biased cross-attention to ligand niche genes No Learns gene embeddings contextualized by ligand signals in spatial microenvironments; aids clustering, ligand–receptor inference, and perturbation simulation 4
@@
FmH2ST wang2025fmh2st Nucleic Acids Research Spatial (Multimodal Foundation Model) Histology image + spatial transcriptomics Image foundation + dual graphs + spot branch No Predict spatial gene expression from histology using fused image and spot features, supporting denoising, heterogeneity detection, and regulatory inference 3
@@
OmiCLIP cui2025towards Nature Methods Spatial (Multimodal Foundation Model) Histology + Spatial Transcriptomics Image–gene contrastive alignment (rank-based) No Learns unified visual-omics embeddings linking histopathology and spatial gene expression; enables cross-modal prediction, annotation, and tissue retrieval 7
@@
QuST-LLM huang2024qust ArXiv Text-Bridge LLM Spatial transcriptomics + histology metadata GO-term + gene enrichment + LLM narrative overlay Yes Converts ST data and ROIs into human-readable narratives and matches natural language queries to spatial regions via GO/LLM interpretation 4
@@
SpaCCC yang2024deciphering IEEE Xplore Text-Bridge LLM Spatial transcriptomics / transcriptomic genes (LRs) LLM embeddings of ligand + receptor genes No Infers spatially resolved cell–cell communication by embedding LR pairs in LLM latent space + diffusion / permutation test filtering 3
@@
spaLLM 10778152 Briefings in Bioinformatics Spatial (Multimodal Foundation Model) Spatial multi-omics (RNA, ATAC, protein) scGPT embeddings + GNN + attention fusion No Enhances spatial domain identification by fusing LLM-derived embeddings and spatial-omics signals via multi-view attention 3
@@
scMMGPT (2025) shi2025multimodal ArXiv Text-Bridge LLM / Multimodal_FM RNA + Text (metadata) Annotation, description No Generation - links single-cell and text PLMs to describe cells, generate pseudo-cells from text, and enhance annotation through text-conditioned reasoning 6
@@
Cell2Sentence (C2S) pmlr-v235-levine24a PMLR Text-Bridge LLM RNA (scRNA-seq converted to “cell sentences”) Marker / Implicit (gene rank-order) No Generation, annotation & reconstruction - encodes cells as gene-ranked sentences and fine-tunes LLMs to classify or generate biologically meaningful cell text 6
@@
C2S-Scale Rizvi2025.04.14.648850 BioRxiv Text-Bridge LLM+ Multimodal Foundation Model RNA + Text / Metadata Atlas + Text / Implicit rank grounding No Chat, generation & annotation - trains large LLMs on cell sentences and biological text to enable perturbation prediction and multicellular summarization 7
@@
Cell2Text kharouiche2025cell2text ArXiv Text-Bridge LLM + Multimodal Foundation Model RNA → Natural Language Gene-level embeddings + ontology / metadata grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict expression and interpret TF–cis interactions 7
@@
GenePT chen2024genept BioRxiv Foundation Model / Text-Augmented scRNA Literature / Text embedding grounding No Embedding & downstream prediction - uses GPT-3.5 gene text embeddings to derive cell embeddings via weighted or ranked gene aggregation. 5-7
@@
CellLM zhao2023large ArXiv Foundation Model / Representation model RNA (single cell expression) Implicit - contrastive embedding from expression data No Represent cells via a contrastive-learning transformer, optimize embedding space for tasks like cell-type annotation, drug sensitivity prediction 6
@@
scExtract wu2025scextract Genome Biology Agentic Framework (Text-Bridge LLM) RNA (scRNA-seq) + Text Article-based parameter extraction Yes Annotation & Integration - uses LLMs to extract pipeline parameters from publications and apply them for dataset harmonization. 7
@@
scAgent mao2025scagent ArXiv Agentic Framework (Text-Bridge hybrid) RNA (scRNA-seq) Reference atlas + marker gene reasoning + memory grounding Yes Universal cell annotation & novel cell discovery - scAgent uses an LLM planning module, memory, and tool modules to annotate cells across tissues, detect unknown types, and incrementally learn new annotations 7
@@
EpiFoundation wu2025epifoundation BioRxiv Epigenomic Foundation Model / Multimodal alignment scATAC (chromatin accessibility) Peak-to-gene supervision (alignment to expression) No Cell embedding, annotation, batch correction, gene expression prediction - trains on sparse peak sets, aligns to gene expression supervision, and transfers learned embeddings for downstream ATAC tasks 7
@@
EpiAgent chen2025epiagent BioRxiv Epigenomic Foundation Model / Multimodal text-bridge hybrid scATAC / chromatin accessibility Peak tokenization + external embeddings + regulatory supervision grounding No Embedding, annotation, imputation, and perturbation prediction - encodes chromatin accessibility as ranked cCRE tokens, enabling zero-shot cell annotation, peak imputation, and response prediction. 8
@@
ChromFound jiao2025chromfound Nature Epigenomic → Expression Transformer / Regulatory FM scATAC (chromatin accessibility) + DNA sequence Motif × peak matrix / masked regulatory grammar grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict gene expression in seen and unseen cell types; also interprets transcription factor interactions and cis-regulatory elements 8
@@
GET (General Expression Transformer) fu2025foundation Nature Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6
@@
scMPT palayew2025towards ArXiv Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6